OpenStudy (anonymous):

Solve the absolute value equation. |x| = 4 |x + 2| = 9 |x – 1| + 3 = 5

OpenStudy (anonymous):

\[\left| x+1 \right|+3=5\] \[\left| x+1 \right|=5-3=2\] \[x+1=\pm2\] either x+1=2 x=2-1=1 or x+1=-2 x=-2-1=-3 similarly you can solve others